OpenStudy (anonymous):

Solve the trigonometric equation sec theta - 2 = 0. Find all solutions in the interval [0, pi]

OpenStudy (anonymous):

let theta=x so, sec x -2 = 0 sec x = 2 cos x = 1/2 cos x = cos pi/3 x = pi/3 i.e theta = pi/3

OpenStudy (anonymous):

Thanks for breaking this down for me. Appreciate it

OpenStudy (anonymous):

welcome